OpenStudy (anonymous):

8 - 6 + 1 / 4^2 + 7 is the answer 3/23 show work

OpenStudy (anonymous):

\[8 - 6 + (1/16) + 7\]?? then 8 - 6 = 2 2 + 7 = 9 9 + 1/16 = (9*16 + 1)/ 16 = 145/16 So, i guess not :P

myininaya (myininaya):

Is the question as written? Or do you mean (8-6+1)/(4^2+7)

myininaya (myininaya):

if you do mean what i wrote, then your answer is right.